OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of 2/3? 2x + 3y = 3 2x – 3y = 3 3x + 2y = 3 3x – 2y = 3

OpenStudy (anonymous):

the line having slope 2/3 is-------- y= 2/3 x + c Since line passes through (6,3).. put x=6 and y =3 3= 2/3 *6 +c 3= 4 +c c= 3-4 = -1 So the equation is-- y=2/3 x -1 or, 3y= 2x-3 or, 2x -3y = 3